OpenStudy (anonymous):

If the entropy change for a system at 100.0°C is -20.0 J/K, then which of the following describes the heat? A. 7.46 kJ of heat added B. 2.00 kJ of heat removed C. 2.00 kJ of heat added D. 7.46 kJ of heat removed *thanks:)

OpenStudy (michele_laino):

hint: this formulas, in general, holds: \[\Large \Delta S = \frac{Q}{T}\]

OpenStudy (michele_laino):

formula*

OpenStudy (anonymous):

-20/100?

OpenStudy (anonymous):

:/

OpenStudy (michele_laino):

no, we have to write the inverse formula, then we have to compute the absolute temperature \[\Large T = t + 273\] \[\Large Q = T\Delta S\]

OpenStudy (anonymous):

ohh okay! so 373 * -20 ? and we get-7460? so that means choice D? 7.46 kJ heat removed? since it is negative?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

congratulations!!!!!

OpenStudy (anonymous):

yayy!! thank you:)

OpenStudy (michele_laino):

thank you!! :D